OpenStudy (anonymous):

how many real number solutions are there to the equation o=-3x^2+x-4???

OpenStudy (anonymous):

a) 0 b) 1 C) 2 D)3

OpenStudy (anonymous):

Are you familiar with the quadratic formula? I suggest you use this to solve the equation and then simply look at how many roots you've got. Once you've done this, you can go back and think about which part of the equation dictated how many results you got. If you have an idea, perhaps test with some other quadratics.

OpenStudy (anonymous):

idk the quadratic formula confuses me so much and i cant figure out how to do it

OpenStudy (anonymous):

The solutions of \[ax^2 + bx + c = 0\]are given by \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

Ik, but i dont understand its really confusing on how to set it up

OpenStudy (anonymous):

What don't you understand? Where it comes from? Otherwise it's just a matter of substituting coefficients in until you're used to it.

OpenStudy (anonymous):

i dont get how to substitute them in

OpenStudy (anonymous):

By comparing the ax²+bx+c with the equation you're trying to solve, you can see the values of the coefficients a, b and c which you can then use in the quadratic formula.

OpenStudy (anonymous):

but there are more than 3 numbers?

OpenStudy (anonymous):

Hm?

OpenStudy (anonymous):

nvm